OpenStudy (anonymous):

Please help :) Find three consecutive odd integers such that the product of the first and the second exceeds the third by eight. the first integer is... the second integer is... the third integer is...

OpenStudy (anonymous):

odd integer means 2n+1 \[(2n+1)(2n+3)+8=2n+5\]

OpenStudy (anonymous):

so whats the first integer?

OpenStudy (anonymous):

hold on i must have messed this up somehow

OpenStudy (anonymous):

not possible

OpenStudy (anonymous):

oh yes it is

OpenStudy (anonymous):

they are 3,5 and 7

OpenStudy (anonymous):

3*5=15

OpenStudy (anonymous):

7+8=15

OpenStudy (anonymous):

i put the stupid 8 on the wrong side!